# Paramagnon-Interference Mechanism for Three-Dimensional Bond Order in Kagome Metals AV3Sb5 (A=Cs, Rb, K): Analysis by the Density-Wave Equation

### v2

@@ III DW Equation Analysis for 3D Kagome Metal Model
Here, we study the BO in the 3D kagome metal models based on the DW equation theory. In this theory, various nonmagnetic DW orders are caused by the beyond-RPA non-local correlations. This theory was first developed to explain the orbital nematic state in Fe-based superconductors, and it has been successfully applied to the cuprates [70] and the twisted-bilayer graphene [71]. Recently, the present authors studied 2D kagome metals in Ref. [14]. By solving the DW equation including the Hartree, Maki-Thompson (MT), and AL terms shown in Fig. 5 (a), we can calculate the infinite series of AL and MT terms. The AL terms represent the PMI mechanism, which gives rise to the BO in kagome metals. The analytic expression of the charge-channel DW equation is
@@
Figure 5: (a) Charge-channel density-wave (DW) eigenvalue equation. The largest eigenvalue $\lambda_{{\bm{Q}}}$ represents the instability of the charge-channel DW at wavevector $\bm{Q}$}}, and the form factor $f$ gives the order parameter. The wavy lines represent the spin susceptibilities. Two AL terms describe the paramagnon interference mechanism that gives various unconventional DW states. (b-c) Obtained $\lambda_{{\bm{Q}}}$ for A=Cs at $n=30.8,31$: (b) $\lambda_{{\bm{Q}}}$ in the $q^{z}=0$ plane, where ${{\bm{q}}}_{K}=\left(\frac{2\pi}{\sqrt{3}},\frac{2\pi}{3}\right)$. (c) Small $q_{1}^{z}$-dependence of $\lambda_{{{\bm{Q}}}_{1}}$ for ${{\bm{Q}}}_{1}=({{\bm{q}}}_{1},q_{1}^{z})$. The BO solution is obtained for any $q_{1}^{z}$. (d) $\lambda_{{{\bm{Q}}}}$ for $\alpha_{S}=0.95,0.98$ in $q^{z}=0$ at $n=31$ of A=Cs. (e) $\alpha_{S}$ dependence of $\lambda_{{{\bm{Q}}}_{1}}$ for $q_{1}^{z}=0$ at $n=31$ of A=Cs.
@@
Here, we explain the obtained numerical results for A=Cs at $T=0.04$ and $U=2.6$, where ${\alpha}_{S}=0.90$. Figure 5 (b) shows the obtained $\lambda_{{\bm{Q}}}$ on the $q^{z}=0$ plane at $n=30.8,31$: $\lambda_{{\bm{Q}}}$ shows the almost commensurate peak at ${{\bm{q}}}={{\bm{q}}}_{1}$, which corresponds to the 2D BO wavevector in kagome metals. Interestingly, the $q_{1}^{z}$-dependence of $\lambda_{{{\bm{Q}}}_{1}}$ is quite small ($\lesssim 1$%), as shown in Fig. 5 (c), reflecting the quasi-2D character of the FSs in kagome metals. We note that the value of $\lambda_{{{\bm{Q}}}_{1}}$ for $q_{1}^{z}=\pi$ is slightly larger than that for $q_{1}^{z}=0$ since the spin susceptibility peaks in $q^{z}\sim\pi$ due to the 3D structure of FSs, which derives the wavevector with $q^{z}\sim\pi$ as discussed in Ref. [31]. In this case, the alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$ is the most favorable, while the TrH (SoD) s-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$ is also favored. Within the DW equation, we cannot distinguish between the TrH s-BO and the SoD s-BO. Figures 5 (d) and (e) show $\alpha_{S}$ dependence of $\lambda_{{{\bm{Q}}}}$ for $q^{z}=0$ at $n=31$, $T=0.04$. $\lambda_{{{\bm{Q}}}_{1}}$ reaches $1$ for $\alpha_{S}\gtrsim 0.97$. Thus, the BO transition temperature $T_{\rm BO}\sim 100$K is obtained in the moderate electron correlation.
@@ IV Discussion
Here, we discuss the 3D structure of CDW in kagome metal based on the obtained results by the 3D DW equation and the 3rd-order GL free energy. For simplicity, we take account of commensurate $q_{m}^{z}=-q_{m}^{z}(=0,\pi)$ BO wavevectors, which correspond to $2\times 2\times 1$ and $2\times 2\times 2$ BO. From the $\bm{Q}$}} dependence of $\lambda_{{\bm{Q}}}$ shown in Figs. 5 and 6, we find that $\lambda_{{{\bm{q}}}_{1},\pi}$ is slightly larger than $\lambda_{{{\bm{q}}}_{1},0}$. For this reason, the commensurate $2\times 2\times 2$ alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$ and nematic TrH (SoD) BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$ are favored based on the 3D DW equation due to the PMI mechanism. On the other hand, the coefficient of the 3rd-order GL term $b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})$ is shown in Fig. 11 (b) in Appendix B. The obtained negative (positive) $b_{1}$ favors the TrH (SoD) BO. We confirmed that $b_{1}(\pi,\pi,0)$ is almost the same as $b_{1}(0,0,0)$ since $q^{z}$ dependence of $b_{1}$ is very small. Thus, $2\times 2\times 2$ TrH (SoD) s-BO competes with $2\times 2\times 1$ TrH (SoD) BO. Combined with the results of the 3D DW equation, the nematic $2\times 2\times 2$ TrH (SoD) s-BO is realized via a first-order transition below $T_{\mathrm{BO}}$ by the 3rd-order GL free energy. In contrast, the contribution from the 3rd-order GL free energy is absent to the alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$, which violates the constraint $q_{1}^{z}+q_{2}^{z}+q_{3}^{z}=0$ (modulo $2\pi$). The alternating v-BO will be realized via a second-order transition if the 3rd-order GL free energy is very small. Based on the $n$ dependence of $b_{1}$, the TrH BO tends to be stabilized in the hole-doped regime, whereas the SoD BO tends to be favored in the electron-doped regime. However, the obtained value of $b_{1}$ is sensitive to $T$. Quantitative evaluation of $b_{1}$ is an important future issue.
@@ Acknowledgements.
We derive the 30 orbital tight-binding model with 15 V $d$-orbitals and 15 Sb $p$-orbitals by using the WIEN2k [65] and Wannier90 [66] software. Figure 9 (a) shows the bandstructures for $\Delta E_{p}=-0.2,0$. The bandstructure around the $\Gamma$ point is composed of Sb $p$-orbitals. By introducing $\Delta E_{p}=-0.2$, energy at the $\Gamma$ point $E_{\Gamma}$ decreases from 0.4-0.4eV to 0.55-0.55eV consistently with ARPES measurements [48, 67]. Figure 9 (b) shows the 3D FS for $\Delta E_{p}=-0.2$ at $n=31$. Figures 10 (a) and (b) show the bandstructures of A=Rb and A=K for $\Delta E_{p}=-0.2$, respectively. In these cases, $E_{\Gamma}\sim-0.6$ is also consistent with ARPES measurements [48].
@@
To understand the 3D $3Q$ BO state, we also study the 3rd-order GL free energy $F^{(3)}=b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})\phi_{1}\phi_{2}\phi_{3}$ in the 3D kagome metal models. ${\bm{\phi}}=(\phi_{1},\phi_{2},\phi_{3})$ is the order parameters, and the potential of the BO state at ${{\bm{Q}}}={{\bm{Q}}}_{m}$ is given as $\phi_{m}{\hat{g}}_{{{\bm{Q}}}_{m}}(k)$. Here, we fix the 2D BO wavevectors ${{\bm{q}}}_{m}$ while their $q_{z}$ components $q_{m}^{z}$ are variables, under the constraint $q_{1}^{z}+q_{2}^{z}+q_{3}^{z}=0$ (modulo $2\pi$). Here, we take account of commensurate $q_{m}^{z}=-q_{m}^{z}(=0,\pi)$ BO wavevectors. Diagrammatic expression of $b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})$ is given in Fig. 11 (a), whose analytical expression is introduced in Ref. [15].

### v3

@@ III DW Equation Analysis for 3D Kagome Metal Model
Here, we study the BO in the 3D kagome metal models based on the DW equation theory. In this theory, various nonmagnetic DW orders are caused by the beyond-RPA non-local correlations. This theory was first developed to explain the orbital nematic state in Fe-based superconductors, and it has been successfully applied to the cuprates [70] and the twisted-bilayer graphene [71]. Recently, the present authors studied 2D kagome metals in Ref. [14]. By solving the DW equation, including the Hartree, Maki-Thompson (MT), and AL terms shown in Fig. 5 (a), we can calculate the infinite series of AL and MT terms. The AL terms represent the PMI mechanism, which gives rise to the BO in kagome metals. The analytic expression of the charge-channel DW equation is
@@
Figure 5: (a) Charge-channel density-wave (DW) eigenvalue equation. The largest eigenvalue $\lambda_{{\bm{Q}}}$ represents the instability of the charge-channel DW at wavevector $\bm{Q}$}}, and the form factor $f$ gives the order parameter. The wavy lines represent the spin susceptibilities. Two AL terms describe the paramagnon interference mechanism that gives various unconventional DW states. (b-c) Obtained $\lambda_{{\bm{Q}}}$ for A=Cs at $n=30.8,31,31.2$: (b) $\lambda_{{\bm{Q}}}$ in the $q^{z}=0$ plane, where ${{\bm{q}}}_{K}=\left(\frac{2\pi}{\sqrt{3}},\frac{2\pi}{3}\right)$. (c) Small $q_{1}^{z}$-dependence of $\lambda_{{{\bm{Q}}}_{1}}$ for ${{\bm{Q}}}_{1}=({{\bm{q}}}_{1},q_{1}^{z})$. The BO solution is obtained for any $q_{1}^{z}$. (d) $\lambda_{{{\bm{Q}}}}$ for $\alpha_{S}=0.95,0.98$ in $q^{z}=0$ at $n=31$ of A=Cs. (e) $\alpha_{S}$ dependence of $\lambda_{{{\bm{Q}}}_{1}}$ for $q_{1}^{z}=0$ at $n=31$ of A=Cs.
@@
Here, we explain the obtained numerical results for A=Cs at $T=0.04$ and $U=2.6$, where ${\alpha}_{S}=0.90$. Figure 5 (b) shows the obtained $\lambda_{{\bm{Q}}}$ on the $q^{z}=0$ plane at $n=30.8,31,31.2$: $\lambda_{{\bm{Q}}}$ shows the almost commensurate peak at ${{\bm{q}}}={{\bm{q}}}_{1}$, which corresponds to the 2D BO wavevector in kagome metals. Interestingly, the $q_{1}^{z}$-dependence of $\lambda_{{{\bm{Q}}}_{1}}$ is quite small ($\lesssim 1$%), as shown in Fig. 5 (c), reflecting the quasi-2D character of the FSs in kagome metals. We note that the value of $\lambda_{{{\bm{Q}}}_{1}}$ for $q_{1}^{z}=\pi$ is slightly larger than that for $q_{1}^{z}=0$ since the spin susceptibility peaks in $q^{z}\sim\pi$ due to the 3D structure of FSs, which derives the wavevector with $q^{z}\sim\pi$ as discussed in Ref. [31]. In this case, the alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$ is the most favorable, while the TrH (SoD) s-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$ is also favored. Within the DW equation, we cannot distinguish between the TrH s-BO and the SoD s-BO. The band structures at $n=30.8$ and $31.2$ are shown in Figs. 11(a) and (b) in Appendix A, respectively. The energy at the $b_{3g}$-orbital vHS $(E_{\rm vHS})$ is obtained as $E_{\rm vHS}=-0.013(0.079)$ at $n=30.8(31.2)$. At $n=31.2$, $T<|E_{\rm vHS}-\mu|$ is satisfied. Even in the condition, a nearly $2\times 2\times 2$ BO is reproduced. Figures 5 (d) and (e) show $\alpha_{S}$ dependence of $\lambda_{{{\bm{Q}}}}$ for $q^{z}=0$ at $n=31$, $T=0.04$. $\lambda_{{{\bm{Q}}}_{1}}$ reaches $1$ for $\alpha_{S}\gtrsim 0.97$. Thus, the BO transition temperature $T_{\rm BO}\sim 100$K is obtained in the moderate electron correlation.
@@
It is noteworthy that the total BO eigenvalue is enhanced as $\lambda_{\bm{q}}^{\rm tot}=\lambda_{\bm{q}}+\gamma_{\bm{q}}$, where $\gamma_{\bm{q}}(>0)$ originates from the electron–phonon interaction, as discussed for Fe-based superconductors [72]. Therefore, even when $\lambda_{\bm{q}}$ is slightly incommensurate (for $n=31.2$), $\lambda_{\bm{q}}^{\rm tot}$ can become commensurate due to the $\bm{q}$-dependence of $\gamma_{\bm{q}}$. Furthermore, the $3Q$ BO transition in kagome metals is a weak first order due to a symmetry-allowed third-order term in the GL free energy. As a result, even if $\lambda_{\bm{q}}$ is slightly incommensurate, a commensurate $2\times 2\times 2$ BO may emerge through a first-order transition driven by lattice deformation. It is expected that the $2\times 2\times 2$ BO in kagome metals arises from the cooperative effects of Coulomb repulsion and electron–phonon interactions.
@@ IV Discussion
Here, we discuss the 3D structure of CDW in kagome metal based on the obtained results by the 3D DW equation and the 3rd-order GL free energy. For simplicity, we take account of commensurate $q_{m}^{z}=-q_{m}^{z}(=0,\pi)$ BO wavevectors, which correspond to $2\times 2\times 1$ and $2\times 2\times 2$ BO. From the $\bm{Q}$}} dependence of $\lambda_{{\bm{Q}}}$ shown in Figs. 5 and 6, we find that $\lambda_{{{\bm{q}}}_{1},\pi}$ is slightly larger than $\lambda_{{{\bm{q}}}_{1},0}$. For this reason, the commensurate $2\times 2\times 2$ alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$ and nematic TrH (SoD) BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$ are favored based on the 3D DW equation due to the PMI mechanism. On the other hand, the coefficient of the 3rd-order GL term $b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})$ is shown in Fig. 12 (b) in Appendix B. The obtained negative (positive) $b_{1}$ favors the TrH (SoD) BO. We confirmed that $b_{1}(\pi,\pi,0)$ is almost the same as $b_{1}(0,0,0)$ since $q^{z}$ dependence of $b_{1}$ is very small. Thus, $2\times 2\times 2$ TrH (SoD) s-BO competes with $2\times 2\times 1$ TrH (SoD) BO. Combined with the results of the 3D DW equation, the nematic $2\times 2\times 2$ TrH (SoD) s-BO is realized via a first-order transition below $T_{\mathrm{BO}}$ by the 3rd-order GL free energy. In contrast, the contribution from the 3rd-order GL free energy is absent to the alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$, which violates the constraint $q_{1}^{z}+q_{2}^{z}+q_{3}^{z}=0$ (modulo $2\pi$). The alternating v-BO will be realized via a second-order transition if the 3rd-order GL free energy is very small. Based on the $n$ dependence of $b_{1}$, the TrH BO tends to be stabilized in the hole-doped regime, whereas the SoD BO tends to be favored in the electron-doped regime. We summarize these results for A=Cs in Table 1. However, the obtained value of $b_{1}$ is sensitive to $T$. Quantitative evaluation of $b_{1}$ is an important future issue.
A=Cs Hole-doped Non-doped Electron-doped
(i) $\lambda$ in DW Eq. v-BO $\gtrsim$ s-BO
(ii) 3rd-order GL $b_{1}<0$ $b_{1}\sim 0$ $b_{1}>0$
(i) + (ii) TrH s-BO v-BO SoD s-BO
Table 1: Summary of the results for A=Cs. By combining the results of (i) $\lambda$ in the DW Eq. and (ii) 3rd-order GL term $b_{1}$, the TrH s-BO, the v-BO, and the SoD s-BO are favored in the hole-doped, non-doped, and electron-doped cases, respectively.
@@ Acknowledgements.
We derive the 30 orbital tight-binding model with 15 V $d$-orbitals and 15 Sb $p$-orbitals by using the WIEN2k [65] and Wannier90 [66] software. Figure 9 (a) shows the band structures for $\Delta E_{p}=-0.2,0$. The band structure around the $\Gamma$ point is composed of Sb $p$-orbitals. By introducing $\Delta E_{p}=-0.2$, energy at the $\Gamma$ point $E_{\Gamma}$ decreases from 0.4-0.4eV to 0.55-0.55eV consistently with ARPES measurements [48, 67]. Figure 9 (b) shows the 3D FS for $\Delta E_{p}=-0.2$ at $n=31$. Figures 10 (a) and (b) show the band structures of A=Rb and A=K for $\Delta E_{p}=-0.2$, respectively. In these cases, $E_{\Gamma}\sim-0.6$ is also consistent with ARPES measurements [48]. Figures 11(a) and (b) show the band structures of A=Cs for $\Delta E_{p}=-0.2$ at $n=30.8$ and $n=31.2$, respectively. The energy at the $b_{3g}$-orbital vHS $E_{\rm vHS}$ is obtained as $E_{\rm vHS}=-0.013(0.079)$ at $n=30.8(31.2)$.
@@
Figure 11: Band structures of A=Cs for $\Delta E_{P}=-0.2$ at (a) $n=30.8$ and (b) $n=31.2$. Energy at the $b_{3g}$-orbital vHS $E_{\rm vHS}=-0.013(0.079)$ at $n=30.8(31.2)$.
@@
To understand the 3D $3Q$ BO state, we also study the 3rd-order GL free energy $F^{(3)}=b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})\phi_{1}\phi_{2}\phi_{3}$ in the 3D kagome metal models. ${\bm{\phi}}=(\phi_{1},\phi_{2},\phi_{3})$ are the order parameters, and the potential of the BO state at ${{\bm{Q}}}={{\bm{Q}}}_{m}$ is given as $\phi_{m}{\hat{g}}_{{{\bm{Q}}}_{m}}(k)$. Here, we fix the 2D BO wavevectors ${{\bm{q}}}_{m}$ while their $q_{z}$ components $q_{m}^{z}$ are variables, under the constraint $q_{1}^{z}+q_{2}^{z}+q_{3}^{z}=0$ (modulo $2\pi$). Here, we take account of commensurate $q_{m}^{z}=-q_{m}^{z}(=0,\pi)$ BO wavevectors. Diagrammatic expression of $b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})$ is given in Fig. 12 (a), whose analytical expression is introduced in Ref. [15].